### v1

# Light-induced torque in ferromagnetic metals via orbital angular momentum generated by photon-helicity

Abstract
We investigated photon-helicity-induced magnetization precession in Co1-xPtx alloy thin films. In addition to field-like torque, attributable to magnetic field generation owing to the inverse Faraday effect, we observed non-trivial and large damping-like torque which has never been discussed for single ferromagnetic layer. The composition dependence of those two torques is effectively elucidated by a model that considers mutual coupling via spin-orbit interaction between magnetization and the electronic orbital angular momentum generated by photon-helicity. This work significantly enhances our understanding of the physics relevant to the interplay of photon-helicity and magnetization in magnetic metals.
Photon carries both spin and orbital angular momenta, in circularly-polarized and vortex electromagnetic wave, respectively, and utilization of those angular momentum of photon is one issue in modern physics[1, 2]. Considerable attention has been devoted in condensed matter physics to the interaction between photon’s angular momentum and matter, fostering technological advances such as vortex-nano-processing[3, 4] and optical information communications[5, 6, 7].
Intriguing physical phenomena relevant to the angular momentum of photon is also discussed as one of the central subjects in opto-magnetism. The inverse Faraday effect is a physical phenomenon in which magnetization ${\bf M}_{\rm p}$ emerges during the dwell time when circularly-polarized light passes through a dielectric medium, and it can be expressed as,
${\bf M}_{\rm p}\propto{\bf E}(\omega_{\rm p})\times{\bf E}^{\star}(\omega_{\rm p}),$ (1)
where ${\bf E}$ is electric field vector for light with an angular frequency of $\omega_{\rm p}$[8, 9]. This instantaneous magnetization can be treated as an effective magnetic field in magnetic dielectric materials. Consequently, GHz-to-THz dynamics of magnetic order are initiated when a circularly-polarized femtosecond laser pulse passes through magnetic materials via the inverse Faraday effect, enabling us to explore various spin dynamics along with various probe, which is an ultrafast and versatile modern technique established to date[10, 11, 12, 13, 14, 15].
The inverse Faraday effect has also been discussed in magnetic metals and a highlight is circularly-polarized laser pulse-induced magnetization reversal, commonly referred to as All-Optical Helicity Dependent Switching (AO-HDS), partially attributed to the inverse Faraday effect[16, 17, 18, 19, 20, 21, 22]. Considerable efforts have been devoted to understanding the physics of the inverse Faraday effect in metals; however, there is still a notable gap in our understanding regarding of the key physics: the nature of magnetization arising from the inverse Faraday effect in metallic ferromagnets. A recent study discussed the inverse Faraday effect in terms of the above-mentioned effective magnetic field to interpret the light-induced precessional magnetization dynamics observed in various ferromagnets and normal metal bilayers[13, 23, 24, 25, 26, 27]. Many theories reported light-induced magnetization of metals[28, 29, 30, 31, 32, 33, 34, 35, 36]. According to the recent theoretical analysis of the inverse Faraday effect magnetization ${\bf M}_{\rm p}$ in metals predominantly consists of orbital angular momentum (OAM) and not spin angular momentum (SAM)[28]. On the other hand, in addition to the inverse Faraday effect, both OAM and SAM can be induced as an absorbed quantity[34, 37, 38, 36]. Although the inverse Faraday effect is proportional to the laser intensity, i.e., the effect vanishes after the laser passes away from the matter, OAM and SAM are induced by the absorption that remains in matter even after the laser passes away from it. By using the absorption, OAM can be induced in any dissipative materials. It is then intriguing to determine whether the light-induced angular momentum in metals is relevant to orbital magnetism in the context of recent research on the magnetization dynamics induced by an electric field (current) via OAM[39, 40, 41, 42, 43, 44, 45]. Furthermore, their deep understanding will also push technologies relevant to opto-magnetism, such as the aforementioned AO-HDS.
Figure 1: (a) and (c) Schematic illustration of photon-helicity generated OAM of electron and magnetization dynamics excitation. Circularly-polarized light has an angular momentum where its sign depends on helicity ($\sigma$). The OAM ($\delta{\bf L}$) is generated parallel to ${\bf S}_{\rm p}$. Magnetization dynamics are excited by SOC. Schematic of angular momentum flow induced by photon-helicity when (b) relaxation to the environment is stronger than SOC and when (d) SOC is stronger than relaxation to environment. When angular momentum flow through SOC dominates over relaxation to environment, $\delta{\bf L}$ rotates within laser pulse duration; then, magnetization dynamics can be excited by OAM transfer torque as shown in Fig. (c).
In this Letter, we tend to elucidate the role of electronic OAM in the dynamics of magnetization induced by circularly-polarized laser pulses in ferromagnets. Our experimental data clearly demonstrates the existence of not only the effective magnetic field, traditionally interpreted as the inverse Faraday effect, but also a non-trivial and large effective magnetic field which is orthogonal to that induced by the inverse Faraday effect. In other words, these correspond to the field- and damping-like torques, respectively. In the former scenario circularly-polarized laser pulse acts as an effective magnetic field which induces field-like torque. On the other hand, the latter scenario where damping-like torque is generated cannot be explained by the traditional inverse Faraday effect alone. Both torques can be understood with a simple model provided, on an equal footing, considering the non-equilibrium OAM induced by the angular momentum of photon and its coupling with magnetization via spin-orbit interactions in ferromagnetic metals.
As mentioned above, we focus on both kinds of torque (or effective magnetic field), and these are illustrated in Fig. 1. In Fig. 1 the macroscopic magnetization of a ferromagnetic metal film is in the film plane (${\bf m}\parallel{\bf x}$) and circularly-polarized laser pulse with ${\bf k}$ vector incidents along the film normal (${\bf k}\parallel{\bf z}$). In this geometry, magnetization precession is induced and the initial phase is determined by the direction of the total torque. The traditional inverse Faraday effect induces ${\bf M}_{\rm p}$ parallel to ${\bf z}$, which will then act on magnetization as an effective magnetic field, i.e., the field-like torque, ${\bf T}\propto{\bf m}\times{\bf M}_{\rm p}$ [Fig. 1(a)], as reported[13, 23, 26]. In the opposite case, the direction of the ${\bf M}_{\rm p}$ induced within a laser pulse duration changes owing to strong coupling between ${\bf M}_{\rm p}$ and ${\bf m}$, leading to the emergence of the damping-like torque, i.e., ${\bf T}\propto{\bf m}\times({\bf M}_{\rm p}\times{\bf m})$ [Fig. 1(c)]. Hereafter, we assume that ${\bf M}_{\rm p}$ is dominantly caused by the electronic OAM $\delta{\bf L}$ induced by light, as shown in Fig. 1. The path of the angular momentum flow is schematically shown in Figs. 1(b) and (d). When relaxation to environment for induced OAM is stronger than the coupling with magnetization, light acts as the field-like torque [Figs. 1(a) and (b)]. However, coupling between induced OAM and magnetization is stronger than relaxation to environment, the damping-like torque as an OAM transfer torque is induced by light irradiation [Figs. 1(c) and (d)]. Consequently, the two cases are characterized by the coupling strength between ${\bf m}$ and $\delta{\bf L}$ and relaxation of $\delta{\bf L}$. In the realm of torque physics induced by the OAM generated by electric current or field, one can intuitively infer that the coupling strength between ${\bf m}$ and $\delta{\bf L}$ is governed by SOC, given that magnetization stems from SAM in transition metal ferromagnets. Hence, the field- and damping-like torques can be tuned by the heavy-metal addition into ferromagnetic metals.
Figure 2: (a) and (b) Circularly-polarized laser pulse induced change in normalized Kerr rotation angle ($\Delta\theta_{\rm K}(t)/\theta_{\rm K}$) in Co and Co35Pt65 alloy films with different photon-helicity using LCP and RCP laser pulses. (c) and (d) Helicity-dependent change in normalized Kerr angle rotation ($(\Delta\theta_{\rm K}({\rm LCP})-\Delta\theta_{\rm K}({\rm RCP}))/(2\theta_{\rm K})$) as a function of time for Co and Co35Pt65 alloy films extracted from (a) and (b). Broken and solid curves are results fitted using a sinusoidal decayed function [Eq. (2)].
Five-nanometer-thick Co1-xPtx alloy thin film samples with different compositions ($x$) were prepared by magnetron sputtering. The stacking structure of the sample is as follows: Si/SiO2 sub./ Co1-xPtx (5) / Al (3) (thickness is in nanometers). The wavelength and pulse duration used for the circularly-polarized laser induced magnetization dynamics measurements were 800 nm and 100 fs, respectively. The pump fluence ($F_{\rm p}$) used for the measurement was fixed at 4.6 J/m2. An in-plane external magnetic field of 2 T was applied during the measurement. Figure 2 shows the effect of left circularly-polarized (LCP) and right circularly-polarized (RCP) laser on normalized Kerr rotation angle $\Delta\theta_{\rm K}\textfractionsolidus\theta_{\rm K}$ to determine the impact on magnetization dynamics in Co [Fig. 2(a)] and Co35Pt65 alloy [Fig. 2(b)]. Magnetization precession was excited after irradiation with a femtosecond laser pulse. The oscillation phase reversed when the photon-helicity was reversed, indicating a photon-helicity induced torque on magnetization. The slightly different amplitudes between the irradiation of the LCP and RCP resulted from thermally induced magnetization precession, likely caused by a slight misalignment between the sample and the applied field. To extract the photon-helicity induced signal, the normalized difference signal ($(\Delta\theta_{\rm K}({\rm RCP})-\theta_{\rm K}({\rm LCP}))\textfractionsolidus(2\theta_{\rm K})$) was taken as shown in Figs. 2(c) and 2(d). This photon-helicity-induced normalized change in the Kerr rotation angle was then fitted by the following equation:
$f(t)=$
$\hskip 5.69046pt\left(\delta m_{\rm z,sin}\sin(2\pi f_{0}t)+\delta m_{\rm z,cos}\cos(2\pi f_{0}t)\right)\exp(-t/\tau_{0}),$ (2)
where $\delta m_{\rm z,sin(cos)}$, $f_{0}$, and $\tau_{0}$ are sine (cosine) amplitude, frequency, and life-time of magnetization precession, respectively. The broken and solid curves in Figs. 2(c) and 2(d) represent the fitted results based on Eq. (2). In the case of Co, the magnetization precession excited by the photon-helicity was sine-like, indicating that the field-like torque was dominant. In contrast, a phase shift and cosine-like components, which are signatures of the damping-like torque, were exhibited by the Co35Pt65 alloy.
Figure 3: (a) Photon-helicity induced magnetization precession amplitude ($\delta m_{\rm sin}$ and $\delta m_{\rm cos}$) as a function of the Pt composition ($x$) in Co1-xPtx alloy. Square and circle symbols are sine and cosine amplitude, respectively. (b) Phase of magnetization precession ($\varphi$) as a function of $x$ extracted from (a).
Figure 3(a) shows $\delta m_{\rm z,sin}$ and $\delta m_{\rm z,cos}$ plotted as functions of the Pt composition ($x$). Both values increased with increasing $x$. In particular, a relatively significant increase of $\delta m_{\rm z,cos}$ was observed with an increase in $x$, compared with that of $\delta m_{\rm z,sin}$. The relative magnitudes of two values significantly changed, resulting in a change in the initial phase $\varphi$ (= $\tan^{-1}(\delta m_{\rm z,cos}/\delta m_{\rm z,sin})$) [Fig. 3(b)]. This phase change indicates that the damping-like torque enhances with increasing Pt concentration in Co1-xPtx, and therefore the SOC correlates with the above-mentioned physical scenario relevant to the OAM.
For a more quantitative discussion of both amplitudes ($\delta m_{\rm z}$) and phase change ($\varphi$), we consider the physics based on the simple dynamical model of magnetization (${\bf m}$) and the induced OAM ($\delta{\bf L}$). The dynamics of $\delta{\bf L}$ can be expressed as Bloch-like equation analogous to that for spin[46]:
$\frac{{\rm d}\delta{\bf L}}{{\rm d}t}={\bf Q}_{\rm L}+\xi^{\prime}\delta{\bf L}\times{\bf m}-\frac{\delta{\bf L}}{\tau},$ (3)
where ${\bf Q}_{\rm L}$ is the generation rate of the electronic OAM and its sign depends on the photon-helicity [Eq. (1)]. Note that the form of ${\bf Q}_{\rm L}$ in Eq. (3) is based on absorption effect, i.e., if there is no relaxation term OAM is preserved even after laser pulse passes away from the matter. As discussed in Fig. 1, we assume the relaxation time ($\tau$) for the non-equilibrium OAM created and introduce SOC strength $\xi^{\prime}$ as the frequency dimension for the coupling between magnetization ${\bf m}$ and $\delta{\bf L}$. Under the steady state condition ${\rm d}(\delta{\bf L})\textfractionsolidus{\rm d}t$ = 0, we obtain the following expression:
$\delta{\bf L}$ $=\frac{\tau}{(\xi^{\prime}\tau)^{2}+1}{\bf Q}_{\rm L}-\frac{\xi^{\prime}\tau^{2}}{(\xi^{\prime}\tau)^{2}+1}{\bf Q}_{\rm L}\times{\bf m}$
$\hskip 28.45274pt+\frac{(\xi^{\prime}\tau)^{2}\tau}{(\xi^{\prime}\tau)^{2}+1}({\bf Q_{\rm L}}\cdot{\bf m}){\bf m}.$ (4)
The low-frequency dynamics of magnetization in ferromagnetic metals can be expressed using the Landau-Lifshitz equation with torque ${\bf T}$ as follows:
$\frac{{\rm d}{\bf m}}{{\rm d}t}=-\gamma\mu_{0}{\bf m}\times{\bf H}_{\rm eff}+{\bf T},$ (5)
where, $\gamma$, $\mu_{0}$, and ${\bf H}_{\rm eff}$ are the gyromagnetic ratio, vacuum permeability, and effective magnetic field, respectively. Here, we assume the torque is due to the SOC ${\bf T}=-\frac{\gamma\xi^{\prime}}{\mu_{\rm at}}{\bf m}\times\delta{\bf L}$, where $\mu_{\rm at}$ is the atomic magnetic moment. Then, we derived the following dynamical equation:
$\frac{{\rm d}{\bf m}}{{\rm d}t}$ $=-\gamma\mu_{0}{\bf m}\times{\bf H}_{\rm eff}+\frac{\gamma}{\mu_{\rm at}}\frac{\xi^{\prime}\tau}{(\xi^{\prime}\tau)^{2}+1}{\bf m}\times{\bf Q}_{\rm L}$
$\hskip 5.69046pt-\frac{\gamma}{\mu_{\rm at}}\frac{(\xi^{\prime}\tau)^{2}}{(\xi^{\prime}\tau)^{2}+1}{\bf m}\times({\bf Q}_{\rm L}\times{\bf m}).$ (6)
The second and third terms are the field- and damping-like torques, respectively, and both are governed by the time scale of $\xi^{\prime}$ and $\tau$, as discussed in Fig. 1. Here, laser was irradiated from the film normal; thus, ${\bf Q}_{\rm L}$ = -$Q_{\rm L}{\bf z}$. Then, the field- and damping-like torques were along ${\bf y}$ and ${\bf z}$ directions. The magnetization precession amplitudes can be derived as,
$\delta m_{\rm z,sin}=\frac{\gamma}{\mu_{\rm at}}\frac{\xi^{\prime}\tau}{(\xi^{\prime}\tau)^{2}+1}\beta Q_{\rm L}\Delta t,$ (7)
$\delta m_{\rm z,cos}=\frac{\gamma}{\mu_{\rm at}}\frac{(\xi^{\prime}\tau)^{2}}{(\xi^{\prime}\tau)^{2}+1}Q_{\rm L}\Delta t,$ (8)
where $\Delta t$ is the laser pulse duration. Here, we use ${\bf H}_{\rm eff}=H{\bf x}-M_{\rm eff}({\bf m}\cdot{\bf z}){\bf z}$ with an external magnetic field ($H$) and effective demagnetizing field ($M_{\rm eff}$). The factor $\beta=\sqrt{H/(H+M_{\rm eff})}$ is related to the elliptical shape of magnetization precession due to the demagnetizing field in the thin film.
Figure 4: (a) Magnetization precession amplitude ($\mu_{\rm at}\delta m_{\rm z,sin}/\beta$ and $\mu_{\rm at}\delta m_{\rm z,cos}$) and (b) phase ($\varphi^{\prime}=\tan^{-1}(\beta\delta m_{\rm z,cos})/\delta m_{\rm z,sin}$) plotted as a function of $\xi^{\prime}\tau$ in the model of photon-helicity induced OAM transfer torque. Broken and solid curve in (a) are results calculated based on $M_{\rm p}\xi^{\prime}\tau\textfractionsolidus(1+(\xi^{\prime}\tau)^{2})$ and $M_{\rm p}(\xi^{\prime}\tau)^{2}\textfractionsolidus(1+(\xi^{\prime}\tau)^{2})$, respectively, where $M_{\rm p}$ corresponds to the orbital magnetization generated by light. Solid curve in (b) is the result calculated based on $\varphi^{\prime}=\tan^{-1}(\xi^{\prime}\tau)$. Here, $\xi$ scales with $Z_{\rm ave}^{2}$ and orbital relaxation time $\tau$ of 1 fs.
The experimental and theoretical amplitudes are compared in Fig. 4(a). In this figure, we plot the normalized values of experimentally-obtained magnetization precession amplitudes ($\mu_{\rm at}\delta m_{\rm z,sin}\textfractionsolidus\beta$ and $\mu_{\rm at}\delta m_{\rm z,cos}$) as functions of $\xi^{\prime}\tau$. Here, we evaluated $\xi^{\prime}$ for Co-Pt alloy with different compositions ($x$) using the values of $\xi^{\prime}$ interpolation of the SOC strength energy for Co and Pt (Co: 0.085 eV, Pt: 0.72 eV)[47] with the well-known scaling rule $\sim Z_{\rm avg}^{2}$[48], where $Z_{\rm avg}$ is the average atomic number of Co-Pt alloy. We also set approximately used fixed value of $\tau$ $\sim$ 1 fs, which approximates photo-exited electron relaxation time, i.e., $\tau$ $\sim$ 1 fs for Co[49]. (more detailed discussion for spin-orbit parameters and the relaxation time for the alloys in Supplemental Material[46]). The theoretical curves in Eqs. (7) and (8) are shown with the broken and solid curves, respectively, in Fig. 4(a) under the assumption of a fixed value of $M_{\rm p}$ = $\gamma Q_{\rm L}\Delta t$ $\sim$ 0.8×10-3 [$\mu_{\rm B}$]. The theoretical magnetization precession phase is given by the ratio between the field- and damping-like torques in Eq. (6):
$\varphi_{\rm the}=\tan^{-1}(\xi^{\prime}\tau).$ (9)
Figure 4(b) shows the theoretical and experimental phases of magnetization precession, i.e., $\varphi_{\rm exp}=\tan^{-1}(\beta\delta m_{\rm z,cos}/\delta m_{\rm z,sin})$ from Eqs. (7) and (8) and $\varphi_{\rm the}$ as a function of $\xi^{\prime}\tau$. In those figures, the experimental data for the amplitude and phase are in good agreement with those obtained from the theoretical model with the same footing.
The pertinent physics discussed thus far for helicity-induced magnetization dynamics is worth noting. As mentioned in the introduction, the inverse Faraday effect is traditionally considered to effectively generate magnetic fields in both magnetic metals and dielectrics. Thermodynamic considerations indicate that the effective magnetic field generated by the inverse Faraday effect is proportional to the Faraday or Kerr effect, which is characterized by magneto-optic constant $Q$. However, our sample showed a reduction of the Kerr effect with an increase in the Pt concentration, contrary to the significant increase in helicity-induced torque with increasing Pt concentration[46]. Recent microscopic theories facilitated further understanding of the nature of field- and damping-like torques in metals via SAM directed to $\delta{\bf S}$ and $\delta{\bf S}\times{\bf m}$, respectively, where $\delta{\bf S}$ is SAM induced by the circularly-polarized light[32]. The theoretical computations predicted that field-like torque was significantly enhanced in ferromagnetic metals with large SOC such as FePt, with the damping-like torque being much smaller than the field-like torque. Also, the emergence of the damping-like torque with increasing Pt concentration is difficult to be interpreted as the SAM generation in ferromagnets, i.e., ${\bf M}_{\rm p}\sim\delta{\bf S}$ in Eq. (1), similar to that observed in GaMnAs[50]. When $\delta{\bf S}$ is generated by circularly polarized light irradiation, the time-scale is characterized by the exchange coupling ($J_{\rm ex}$) of the induced $\delta{\bf S}$ and ${\bf m}$, and relaxation time of the induced spin ($\tau_{\rm S}$). Hence, we can expect $\varphi\sim\tan^{-1}(J_{\rm ex}\tau_{\rm S}\textfractionsolidus\hbar)$, similar to Eq. (9). It is difficult to consider that the phase increases with increasing Pt concentration ($x$) in Co-Pt alloys because Pt addition would deteriorate ferromagnetism, that is reduction of $J_{\rm ex}$, and reduce relaxation time of SAM due to its strong SOC. Similarly, the damping-like torque has been discussed in terms of the SAM induced in the heavy-metals in ferromagnet / heavy-metal bilayers; however, the physics behind this torque is related to the spin-transfer-torque owing to the SAM flow in the bilayer[24]. Thus, we ruled out the physical mechanisms of the damping-like torque proposed to date.
Our study indicates that the OAM induced by light plays a much more important role than the SAM induced by the light. This consideration is in good agreement with the fact that the induced OAM is much larger than the spin by a factor of 10, as discussed in the recent theoretical study on the inverse Faraday effect in metals[28]. Furthermore, $M_{\rm p}$ = $\gamma Q_{\rm L}\Delta t$ $\sim$ 0.8 $\times$ 10-3 [$\mu_{\rm B}$] is closer to the order of the magnitude evaluated for the OAM, i.e., $\sim$ 1.3 $\times$ 10-3 [$\mu_{\rm B}$] in Co or $\sim$ 4.0 $\times$ 10-3 [$\mu_{\rm B}$] in Pt at 1.5 eV pump[28, 46] although the above theory of orbital magnetization generated by light is zero-th harmonic effect[51, 38], namely, orbital magnetization is induced only when laser pulse is in the material. Magnetization precession driven by SOC is induced by time-integration of $Q_{\rm L}$ in Eqs. (7) and (8), therefore, we expect that absorption effect has a much stronger contribution than zero-th harmonic effect. We also consider generation of orbital magnetization owing to light absorption in metals[36], which is proportional to the laser fluence. The calculation of the orbital magnetization generated due to light absorption leads to a value one order of magnitude larger than that used to explain the experimental result[46]. A more quantitative discussion requires fully quantum mechanical theory on the OAM induced-torque beyond the naive phenomenological model discussed here, which is out-of-scope of the Letter and will be investigated in future works.
In summary, we investigated helicity-dependent laser-induced magnetization precession in Co1-xPtx alloys. We demonstrated that the magnitude and the phase of the magnetization precession induced by the circularly-polarized laser pulse systematically increase with the Pt concentrations in the Co1-xPtx alloy film. Those evolutions are attributed to two different types of induced torques, field- and damping-like torques. The experimental data are well described by a simple model that considers the electronic OAM generated by light and its coupling with magnetization via SOC. This study deepens the understanding of the physics relevant to the interaction with photon-helicity and matter. Furthermore, this work opens alternative route for manipulating magnetization in magnetic substances.
This work was partially supported by Grants-in-Aid for Scientific Research (Nos. 21H04648 and 21H05000), JST PRESTO (No. JPMJPR22B2), X-NICS of MEXT (No. JPJ011438), CSIS cooperative research project in Tohoku University, the Asahi Glass Foundation, the Murata Science Foundation, Advanced Technology Institute Research Grants, FRIS Creative Interdisciplinary Collaboration Program in Tohoku University. This work was supported by the ANR-20-CE09-0013 UFO, by “Lorraine Université d’Excellence” reference ANR-15-IDEX-04-LUE, and by the French National Research Agency through the France 2030 government grants PEPR Electronic EMCOM (ANR-22-PEEL-0009). S. I. thanks to TI-FRIS fellowship at Tohoku University. K. N. and K. I. thank to GP-Spin at Tohoku University. S. M. thanks to CSRN of CSIS at Tohoku University. The authors thank Gyung-Min Choi for valuable discussion.